OpenStudy (anonymous):

I really need some help with parabolas. like how to convert the vertex form to standard form. Is what I really need help learning right now.

OpenStudy (anonymous):

My vertex equation is -->\[f(x)=2(x-10)^2+7\] @phi

OpenStudy (phi):

in that case, you want to learn how to multiply out (x-10)^2 which means (x-10)(x-10)

OpenStudy (phi):

See if this helps you https://www.khanacademy.org/math/algebra/multiplying-factoring-expression/multiplying-binomials/v/multiplying-binomials what do you get for (x-10)(x-10) ?

OpenStudy (anonymous):

\[x^2-10x-10x-100\] I have to go really quick I'll be back in about 15- 20 minutes

OpenStudy (phi):

ok, looks good so far. notice that -10 x's take away another 10 x's gives -20x so simplify to \[ x^2 -20x -100 \] ( you should expect to see an x^2 term, an x term and a constant (a number) term) now you have \[ f(x)=2(x-10)^2+7 \\ f(x)=2(x^2 -20x -100)+7 \] the next step is distribute the 2 that means multiply 2 times each term inside the parens then simplify to get only 3 terms ( an x^2 term , an x term and a constant term) what do you get ?

OpenStudy (phi):

in other words you should write \[ f(x)=2(x^2 -20x -100)+7 \\ = 2\cdot x^2 + 2\cdot -20 \cdot x + 2 \cdot -100) + 7 \] where I put in dots to show the multiply. But in algebra, we normally leave out the dots after we simplify things. in other words 2*x^2 is written 2x^2 (means the same thing.. just shorter) 2*-20*x is -40*x or -40x 2 * -100 is -200

OpenStudy (anonymous):

okay i am following you. So after I solve i should get --> \[f(x)=2x^2-40x-193\] because you add 7 to the -200

OpenStudy (phi):

perfect. to go back the other way... you "complete the square"

OpenStudy (anonymous):

okay I see. Thank you! You always give me to best help, and I actually learn. I really don't like just getting the answer.

OpenStudy (phi):

as a check, the x coordinate of the vertex when you have the equation in the form \[ f(x) = a x^2 + b x + c \] is -b/(2a) here your b is -40 and a is 2 we get -(-40)/(2*2) = 40/4= 10 that matches the 10 in the vertex form of the equation \[ f(x)=2(x-10)^2+7 \]

OpenStudy (anonymous):

okay, I see and I'll defiantly use that.

OpenStudy (phi):

good luck.

OpenStudy (anonymous):

wait. would c = 193